OpenStudy (elite):

If m = 3 and n = 2, then m 2 - n 2 is equal to _____.

OpenStudy (anonymous):

32-22

OpenStudy (elite):

actualy still need help

pooja195 (pooja195):

@elite Do you mean... \[\huge~\rm~\bf~m^2-n^2 \]

pooja195 (pooja195):

?

OpenStudy (elite):

yes

OpenStudy (mathmale):

I can only guess what you (elite) mean by "m 2 - n 2". If y ou meant exponentiation, please write those expressions as pooja195 has done. Please don't make others guess regarding your intentions.

pooja195 (pooja195):

Sub the given values \[\huge~\rm~\bf~3^2-2^2\]

pooja195 (pooja195):

3^2=3*3=?